# A Risk-Based Equilibrium Analysis of Energy Imbalance Reserve in Day-Ahead Electricity Markets

### v1

@@ I Background
New England’s power system has been relying on natural gas-fired generating units. In fact, 55% of electricity in 2024 was produced from natural gas [isoresourcemix]. Natural gas is primarily delivered to New England via interstate pipelines, making it especially vulnerable during winter cold spells when residential heating demand surges. During cold spells, gas supply is often constrained and New England gas power plants that rely on just-in-time delivery of natural gas face challenges of procuring fuel in real time. These gas plants often do not have sufficient incentives to make advance arrangements with gas suppliers if they do not expect to operate in real time (i.e. do not have any day-ahead market awards). Of all the ISO regions in the US, ISO New England (ISO NE)’s gas power fleet have the second lowest rate of firm contracts, with 45% of plants purchasing firm contracts in 2024 [eia923]. Instead, they opt for interruptible contracts on the spot market, which puts them at the lowest priority for fuel delivery when the natural gas infrastructure is constrained.  [^1] In the models presented in Section II, we assume that generators are able to choose between investing in fuel at a single fixed price in advance of real time generation or purchase fuel with a range of different prices that correlate with real time electricity demand on a spot market. The choice between fixed and uncertain fuel costs are not identical to the choice between interruptible or firm fuel contracts, but serve as a proxy for this investment choice on a shorter time scale.
In addition, there is frequently a gap between the energy cleared in the day-ahead market (DAM) and the real-time load. In 2023, this gap was positive in 64% of all hours in 2023, with an hourly average value of 515.7 MW and a maximum value of 3397.9 MW during hours where the energy cleared in the DAM is less than the real-time load [energygapdataset]. Similar persistent real time and day-ahead imbalances exist in other US ISO/RTO regions, even prompting the California system operator (CAISO) to create a new imbalance reserve product [caiso_ir], [california_iso_day_2023]. Because additional resources needed to meet the gap are often procured after the DAM clears through the Reserve Adequacy Analysis (RAA) process, market clearing prices in the DAM do not reflect the actual cost of meeting real-time load, adversely affecting market transparency. More importantly, participants do not have strong incentives to procure fuel in advance if they do not have any DAM obligations.
@@
In order to address these challenges [white], ISO NE introduced an enhanced DAM on February 28th, 2025, including a novel ancillary service product called (EIR) [clements_186_nodate]. EIR is designed to balance the gap between the real-time load predicted by ISO NE and energy cleared in the DAM. The day-ahead market will now clear both day-ahead energy and energy imbalance reserve to meet the load forecast. EIR is an energy call ”option” product in the DAM where sellers receive an EIR payment in DAM and potentially may incur a close-out cost based on real-time energy prices and a strike price set by the ISO.
@@ II Modeling Framework, Literature Review and Contributions
The model follows a two-stage market structure, consistent with ISO’s practices [isone_ed], [hoganpjmed], [ferc_ed]. In the first stage (day-ahead), participants make decisions based on foretasted real-time conditions. In the second stage (real-time), under each scenario, which represents a possible real-time demand and fuel price realization, only generators make decisions on whether or not to purchase additional fuel and real time production quantities.
@@
To capture the impact of risk preference on market outcomes and participants’ behaviors, the objective functions of participants’ profit maximization problem use Conditional Value-at-Risk (CVAR). CVAR represents the generator’s expected profits that fall below a percentile of the profit distribution. We adopt the CVAR formulation from Rockafellar and Uryasev [cvar].
@@
The resulting market equilibrium is the solution to a mixed complementarity problem (MCP) derived from the optimality conditions of the agents’ profit maximization problems and market-clearing constraints. The use of mixed complementarity problems to model the equilibrium of multiple agents participating in an electricity market is well established in the operations research literature on power markets. [dart_eqb] models generators participating in both day-ahead and real time markets. [gencaprisk] and [smeers_co-optimization_2022] use two-stage stochastic equilibrium models to explore capacity expansion with risk-averse agents and utilize CVAR as their risk measure. [gurkan_generation_nodate] and [de_maere_daertrycke_investment_2017] use MCPs to compare different market structures and policies, specifically how the presence of risk mitigation instruments changes investment behavior and social welfare. A full treatment of mixed complementarity modeling in energy systems can be found in [gabriel_complementarity_2013] and a review of applications in electricity markets can be found in [dimitriadis_review_2021] and [ruiz_tutorial_2014]. Also, see [murphy_tutorial_2016] for an introduction and tutorial on the use of complementarity problems to analyze different market policies. All the MCPs discussed in this paper were implemented in GAMS and solved using the PATH solver [dirkse_path_1995].
@@ III-A1 Generators’ Profit Maximization Problem in EMO
The generator’s objective function captures the CVAR of its total day-ahead and real-time profit across uncertain real-time scenarios. This formulation follows the approach introduced by Rockafellar and Uryasev [cvar]. In the model, the generator maximizes its risk adjusted profit by choosing a benchmark profit level $\eta_{i}$ and penalizing the expected profit below that level using $u_{i,s}$. The non-negative variable $u_{i,s}$ reflects how much the actual profit in scenario $s$ is below this benchmark profit level $\eta_{i}$ via the first inequality constraint. In other words, The generator agent $i$ hedges against its lowest $\alpha_{i}$ quantile of profits. The parameter $\alpha_{i}\in(0,1]$ corresponds to the generator’s level of risk aversion. When $\alpha_{i}=1$, the generator behaves in a risk-neutral manner and maximizes its expected profit. As $\alpha_{i}$ decreases, the generator becomes more risk-averse, emphasizing performance in the tail of the profit distribution. This CVAR formulation allows us to examine generator’s behaviors under different risk preferences while keeping the model tractable.
@@ IV Theoretical Results of the Risk Neutral Case
Suppose that total available capacity is greater than $FER$ (i.e. $\sum_{i}Q_{i}>FER$). Then any valid solution to the model with energy imbalance reserve and virtual traders, satisfies $\rho=0$. Furthermore, when the close out cost is not identically zero across all scenarios, $e_{i}=0$ in equilibrium.
@@ V-B3 Risk Averse Demand Agent
The introduction of EIR is one possible policy choice to incentivize advanced fuel investment and raise day-ahead energy to the level of the system operator’s load forecast. Consider another choice, where the system operator simply constrains the day-ahead generation to be greater than or equal to the load forecast, $\sum_{i}g_{i}^{DA}\geq FER$. Many system operators do in fact schedule units ahead of real time using a load forecast in place of demand bids, in a an out-of-market process that goes by different names across ISO/RTO regions (Reserve Adequacy Analysis in ISO-NE, Reliability Unit Commitment in ERCOT/PJM, Residual Unit Commitment in CAISO, etc. [ecco_international_design_nodate]). In this section, we examine what would occur if the load forecast constraint enforced in the out of market reliability commitment process was brought into the day-ahead market and compare this to the EMIR model results. The full MCP formulation of the energy market with load forecast constraint is presented in Appendix IX.B.

### v2

@@ I Background
New England’s power system has been relying on natural gas-fired generating units. In fact, 55% of electricity in 2024 was produced from natural gas [11]. Natural gas is primarily delivered to New England via interstate pipelines, making it especially vulnerable during winter cold spells when residential heating demand surges. During cold spells, gas supply is often constrained and New England gas power plants that rely on just-in-time delivery of natural gas face challenges of procuring fuel in real time. These gas plants often do not have sufficient incentives to make advance arrangements with gas suppliers if they do not expect to operate in real time (i.e. do not have any day-ahead market awards). Of all the ISO regions in the US, ISO New England (ISO NE)’s gas power fleet have the second lowest rate of firm contracts, with 45% of plants purchasing firm contracts in 2024 [1]. Instead, they opt for interruptible contracts on the spot market, which puts them at the lowest priority for fuel delivery when the natural gas infrastructure is constrained.  [^1] In the models presented in Section II, we assume that generators are able to choose between investing in fuel at a single fixed price in advance of real time generation or purchase fuel with a range of different prices that correlate with real time electricity demand on a spot market. The choice between fixed and uncertain fuel costs are not identical to the choice between interruptible or firm fuel contracts, but serve as a proxy for this investment choice on a shorter time scale.
In addition, there is frequently a gap between the energy cleared in the day-ahead market (DAM) and the real-time load. In 2023, this gap was positive in 64% of all hours in 2023, with an hourly average value of 515.7 MW and a maximum value of 3397.9 MW during hours where the energy cleared in the DAM is less than the real-time load [10]. Similar persistent real time and day-ahead imbalances exist in other US ISO/RTO regions, even prompting the California system operator (CAISO) to create a new imbalance reserve product [21], [15]. Because additional resources needed to meet the gap are often procured after the DAM clears through the Reserve Adequacy Analysis (RAA) process, market clearing prices in the DAM do not reflect the actual cost of meeting real-time load, adversely affecting market transparency. More importantly, participants do not have strong incentives to procure fuel in advance if they do not have any DAM obligations.
@@
In order to address these challenges [22], ISO NE introduced an enhanced DAM on February 28th, 2025, including a novel ancillary service product called (EIR) [3]. EIR is designed to balance the gap between the real-time load predicted by ISO NE and energy cleared in the DAM. The day-ahead market will now clear both day-ahead energy and energy imbalance reserve to meet the load forecast. EIR is an energy call ”option” product in the DAM where sellers receive an EIR payment in DAM and potentially may incur a close-out cost based on real-time energy prices and a strike price set by the ISO.
@@ II Modeling Framework, Literature Review and Contributions
The model follows a two-stage market structure, consistent with ISO’s practices [17], [14], [4]. In the first stage (day-ahead), participants make decisions based on foretasted real-time conditions. In the second stage (real-time), under each scenario, which represents a possible real-time demand and fuel price realization, only generators make decisions on whether or not to purchase additional fuel and real time production quantities.
@@
To capture the impact of risk preference on market outcomes and participants’ behaviors, the objective functions of participants’ profit maximization problem use Conditional Value-at-Risk (CVAR). CVAR represents the generator’s expected profits that fall below a percentile of the profit distribution. We adopt the CVAR formulation from Rockafellar and Uryasev [18].
@@
The resulting market equilibrium is the solution to a mixed complementarity problem (MCP) derived from the optimality conditions of the agents’ profit maximization problems and market-clearing constraints. The use of mixed complementarity problems to model the equilibrium of multiple agents participating in an electricity market is well established in the operations research literature on power markets. [2] models generators participating in both day-ahead and real time markets. [9] and [20] use two-stage stochastic equilibrium models to explore capacity expansion with risk-averse agents and utilize CVAR as their risk measure. [13] and [5] use MCPs to compare different market structures and policies, specifically how the presence of risk mitigation instruments changes investment behavior and social welfare. A full treatment of mixed complementarity modeling in energy systems can be found in [12] and a review of applications in electricity markets can be found in [6] and [19]. Also, see [16] for an introduction and tutorial on the use of complementarity problems to analyze different market policies. All the MCPs discussed in this paper were implemented in GAMS and solved using the PATH solver [7].
@@ III-A1 Generators’ Profit Maximization Problem in EMO
The generator’s objective function captures the CVAR of its total day-ahead and real-time profit across uncertain real-time scenarios. This formulation follows the approach introduced by Rockafellar and Uryasev [18]. In the model, the generator maximizes its risk adjusted profit by choosing a benchmark profit level $\eta_{i}$ and penalizing the expected profit below that level using $u_{i,s}$. The non-negative variable $u_{i,s}$ reflects how much the actual profit in scenario $s$ is below this benchmark profit level $\eta_{i}$ via the first inequality constraint. In other words, The generator agent $i$ hedges against its lowest $\alpha_{i}$ quantile of profits. The parameter $\alpha_{i}\in(0,1]$ corresponds to the generator’s level of risk aversion. When $\alpha_{i}=1$, the generator behaves in a risk-neutral manner and maximizes its expected profit. As $\alpha_{i}$ decreases, the generator becomes more risk-averse, emphasizing performance in the tail of the profit distribution. This CVAR formulation allows us to examine generator’s behaviors under different risk preferences while keeping the model tractable.
@@ IV Theoretical Results of the Risk Neutral Case
Suppose that total available capacity is greater than FERFER (i.e. $\sum_{i}Q_{i}>FER$). Then any valid solution to the model with energy imbalance reserve and virtual traders, satisfies $\rho=0$. Furthermore, when the close out cost is not identically zero across all scenarios, $e_{i}=0$ in equilibrium.
@@ V-B3 Risk Averse Demand Agent
The introduction of EIR is one possible policy choice to incentivize advanced fuel investment and raise day-ahead energy to the level of the system operator’s load forecast. Consider another choice, where the system operator simply constrains the day-ahead generation to be greater than or equal to the load forecast, $\sum_{i}g_{i}^{DA}\geq FER$. Many system operators do in fact schedule units ahead of real time using a load forecast in place of demand bids, in a an out-of-market process that goes by different names across ISO/RTO regions (Reserve Adequacy Analysis in ISO-NE, Reliability Unit Commitment in ERCOT/PJM, Residual Unit Commitment in CAISO, etc. [8]). In this section, we examine what would occur if the load forecast constraint enforced in the out of market reliability commitment process was brought into the day-ahead market and compare this to the EMIR model results. The full MCP formulation of the energy market with load forecast constraint is presented in Appendix IX.B.